OpenStudy (anonymous):

The measures of the angles of a triangle are in the extended ratio 2 : 6 : 7. What is the measure of the smallest angle?

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

let the angles be 2 x ,6x,7x then 2x+6x+7x=180 find x and then the angles.

OpenStudy (anonymous):

24

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

can you help with more @surjithayer

OpenStudy (anonymous):

i will try.

OpenStudy (anonymous):

i think its a

OpenStudy (anonymous):

how can you say ?

OpenStudy (anonymous):

can you please give me the answer this is one of the last ones

OpenStudy (anonymous):

in a right triangle triangle hypotenuse is greater than the sides.

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

\[\tan \theta=\frac{ -270 }{ -480 }=\frac{ 9 }{ 16 },\theta=\tan^{-1} \frac{ 9 }{ 16 }=about ~29 \degree\]

OpenStudy (anonymous):

so you are correct.